# Free energy in spin glass models with conventional order

### v1

@@ Abstract.
Recently, [1] considered spin glass models with additional conventional order parameters characterizing single-replica properties. These parameters are distinct from the standard order parameter used to measure correlations between replicas. A “min-max” formula for the free energy was prescribed in [1]. We rigorously verify this prescription in the setting of vector spin glass models featuring additional deterministic spin interactions. Notably, our results can be viewed as a generalization of the Parisi formula for vector spin glass models in [24], where the order parameter for self-overlap is already present.
@@ 1. Introduction
It is well understood that the classical Sherrington–Kirkpatrick (SK) model has one order parameter that characterizes the correlations between replicas of the system. We call this order parameter the replica order parameter. In more general models, there can be other order parameters. For instance, in the vector spin glass, the self-overlap of a single replica comes into play. We call order parameters that characterize the properties of a single replica the conventional order parameter. In models where both types of order exist, [1] recently proposed a “min-max” prescription for the variational formula of free energy. We rigorously verify this in the setting of vector spin glass.
More precisely, we consider a Hamiltonian of the form $H_{N}(\sigma)+G_{N}(\sigma)$ at size $N$. Here, $H_{N}(\sigma)$ is the standard spin glass Hamiltonian with Gaussian disorder, and $G_{N}(\sigma)$ accounts for additional deterministic spin interactions. Let $\pi$ be the replica order parameter and $m$ be the conventional order parameter in this model. We show that the limit of free energy is given by
$\lim_{N\to\infty}-\frac{1}{N}\mathbb{E}\log\int e^{H_{N}(\sigma)+G_{N}(\sigma)}\mathrm{d}\sigma=\inf_{m}\sup_{\pi}\mathcal{P}(\pi,m)$
for some Parisi-type functional $\mathcal{P}$, verifying [1, (12)]
@@
Our result allows for more general $m$. For instance, $m$ can be the parameter for the mean magnetization, moments of spins, the self-overlap, and combinations of these quantities. Regardless of the choice, $m$ has always to be optimized after the replica order $\pi$.
Our proof is based on interpolating between the free energy with Hamiltonian $H_{N}(\sigma)+G_{N}(\sigma)$ and the free energy with Hamiltonian $H_{N}(\sigma)-\frac{1}{2}\mathbb{E}H_{N}(\sigma)^{2}$ along a Hamilton–Jacobi equation. The simplest form of this technique exists for the Curie–Weiss model (see [10, Chapter 3]). Variants of this have been used to handle the self-overlap order parameter in [18, 7]. We will see that the free energy with the latter Hamiltonian is “pure” in the sense that there is only the replica order in the model, the same as the SK model.
@@ 1.1. Setting
[…]
@@
We call the term $-\frac{N}{2}\xi\left(\frac{\sigma\sigma^{\intercal}}{N}\right)$ the self-overlap correction, which is is equal to $-\frac{1}{2}\mathbb{E}H_{N}(\sigma)^{2}$ and resembles the drift term in an exponential martingale. It has been proved in [8, Corollary 8.3] that
@@
where the Parisi-type functional $\mathscr{P}$ is defined later in (2.3). Here, $\Pi$ is the collection of replica order parameters and there is no conventional order parameter.
@@
The additional term $NG(\cdots)$ is of the type in the generalized Curie–Weiss model. The natural order parameter arising from $NG(\cdots)$ characterizes the limit of $m_{N}$.
@@ 1.2. Main results
The free energy $F^{\mathsf{soc}}_{N}$ in (1.2) with self-overlap is “pure” since there is only the replica order in the formula. Further, we will show in Proposition A.1 of Appendix A that, with the self-overlap correction, all conventional orders are trivial in the sense that $m_{N}$ converges to a constant for any bounded $h$.
@@
In view of Corollary 1.2, without the self-overlap correction, it is necessary to include the self-overlap as part of the conventional order. From (1.12), it is also a minimal requirement. One can interpret this as that the self-overlap is the canonical conventional order parameter. This is not surprising because the self-overlap together with the replica order (which characterizes off-diagonal overlaps) completely describes the entire overlap array.
@@ 1.3. Related works
As mentioned before, the source of motivation is from [1], which used the “min-max” prescription to clarify the relationship between quenched free energies, annealed free energies, and replica order. An earlier work that considered a more restrictive class of spin glass models with conventional order parameters is [14]. Since the self-overlap is one particular conventional order parameter, these connect with the author’s recent studies [7, 6] on the self-overlap in vector spin glasses. The idea of adding the self-overlap correction first appeared in the Hamilton–Jacobi approach to spin glass models [16, 18, 15, 17]. For more detail on this approach, we refer to [10].
Since the theme of this work is to find a formula for free energy, we review works along this direction. Parisi initially proposed the formula for the free energy in the SK model in [25, 26]. Guerra rigorously proved the upper bound in [11] and Talagrand proved the matching lower bound in [30]. Extensions were made by Panchenko to various settings: the SK model with soft spins [19], the scalar mixed $p$-spin model [21, 20], the multi-species model [22], and the mixed $p$-spin model with vector spins [23, 24]. The Parisi formula for the balanced Potts spin glass was recently established in [4]. For spherical spins, Parisi-type variational formulas have been proved for the SK model [29], the mixed $p$-spin model [9], and the multi-species model [3].
@@ 2. Proofs
The goal is to prove our main results. We start with some definitions and conditions on $\xi$. Let $\mathbf{S}^{D}_{+}$ denote the subset of $\mathbf{S}^{D}$ comprising of positive semi-definite matrices. A useful fact (see [12, Theorem 7.5.4]) is that if $a\in\mathbf{S}^{D}$, then
@@
The derivative $\nabla\xi:\mathbb{R}^{D\times D}\to\mathbb{R}^{D\times D}$ is defined with respect to the entry-wise inner product. This encompasses a broad class of models and we refer to [17, Section 6] for examples.
@@ 2.1. Parisi-type functional
[…]
@@
To define the Parisi functional, we recall the Ruelle probability cascade [28]. Let $\mathfrak{R}$ denote the Ruelle probability cascade with overlap uniformly distributed over $[0,1]$ (see [20, Theorem 2.17]). Precisely, $\mathfrak{R}$ is a random probability measure on the unit sphere in a separable Hilbert space, with the inner product denoted by $\alpha\wedge\alpha^{\prime}$. Independently sampling $\alpha$ and $\alpha^{\prime}$ from $\mathfrak{R}$, the law of $\alpha\wedge\alpha^{\prime}$ under $\mathbb{E}\mathfrak{R}^{\otimes 2}$ is the uniform distribution over $[0,1]$, where $\mathbb{E}$ integrates the randomness inherent in $\mathfrak{R}$. Almost surely, the support of $\mathfrak{R}$ is ultrametric in the induced topology. For rigorous definitions and comprehensive properties, we direct the reader to [20, Chapter2] (also see [10, Chapter5]).
@@
For the construction and properties of this process, we refer to [8, Section 4 and Remark 4.9].
@@
Here, $\mathbb{E}$ integrates the Gaussian randomness in $w^{\nabla\xi\circ\pi}(\alpha)$ and then the randomness of $\mathfrak{R}$. It is noteworthy that $-\frac{1}{2}\nabla\xi\circ\pi(1)\cdot\tau\tau^{\intercal}=-\frac{1}{2}\mathbb{E}\left(w^{\nabla\xi\circ\pi}(\alpha)\cdot\tau\right)^{2}$, is the self-overlap correction in this functional.
@@ 2.2. Preliminaries
~$\widetilde{\mathcal{F}}_{N}(t,x)=\frac{1}{N}\log\int\exp\left(H_{N}(\sigma)-\frac{N}{2}\xi\left(\frac{\sigma\sigma^{\intercal}}{N}\right)+tNG\left(m_{N}\right)+Nx\cdot m_{N}\right)\mathrm{d}P_{N}(\sigma).$
@@
The convergence at $x=0$ has been proved in [8, Corollary 8.3] (which is exactly (1.3)). For $x\neq 0$, we can substitute $\mathrm{d}P_{1}$ with $e^{x\cdot h(\tau)}\mathrm{d}P_{1}(\tau)$ and notice that the new measure still satisfies (H0). The convergence at $x$ follows from this substitution. ∎
@@
It is classical that (e.g. [27, Theorem 25.5]) the combination of convexity and differentiability implies continuous differentiability. Hence, it remains to show that $\mathscr{P}^{h}$ is differentiable everywhere. Fix any $x\in\mathbb{R}^{d}$. A vector $a\in\mathbb{R}^{d}$ is said to be a subdifferential of $\mathscr{P}^{h}$ at $x$ if $\mathscr{P}^{h}(y)-\mathscr{P}^{h}(x)\geqslant a\cdot(y-x)$ for every $y\in\mathbb{R}^{d}$. Since $\mathscr{P}^{h}$ is convex, it suffices to show that any subdifferential $a$ of $\mathscr{P}^{h}$ at $x$ is unique. For each $\varepsilon>0$, we choose $\pi_{\varepsilon}$ to satisfy
@@
  • for each $N$, $\mathcal{F}_{N}$ is Lipschitz and convex jointly over +×$\mathbb{R}_{+}\times\mathbb{R}^{d}$;
@@
[…]
@@ 2.3. A PDE approach
If $f$ is a limit of $\mathcal{F}_{N}$, then Lemma 2.1 implies that the relevant initial condition should be $f(0,\cdot)=\mathscr{P}^{h}$, which is continuously differentiable by Lemma 2.2. Also, Lemma 2.3 ensures that $f$ is Lipschitz and convex. These two properties make the following convex selection principle useful in this setting. We refer to [10, Theorem 3.21 and Corollary 3.24] for the proof. This result first appeared in [5].
@@
As $N\to\infty$, $\mathcal{F}_{N}$ converges in the local uniform topology to the unique viscosity solution $f$ of (2.13) with initial condition $f(0,\cdot)=\mathscr{P}^{h}$.
@@
Since $\mathcal{F}_{N}$ is Lipschitz uniformly in $N$ (by Lemma 2.3), the Arzelà–Ascoli theorem implies that any subsequence of $\mathcal{F}_{N}$ has a further subsequence that converges in the local uniform topology to some $f$. By the uniqueness of the solution (e.g. [10, Corollary 3.7]), it suffices to show that any such $f$ is the viscosity solution. For lighter notation, we assume that the entire sequence $\mathcal{F}_{N}$ converges to $f$.
@@
Using the convexity of ~$\widetilde{\mathcal{F}}_{N}$ due to (2.12), we have
~$\widetilde{\mathcal{F}}_{N}(t_{N},x_{N}+y)\geqslant\widetilde{\mathcal{F}}_{N}(t_{N},x_{N})+y\cdot\nabla\widetilde{\mathcal{F}}_{N}(t_{N},x_{N}).$
@@
By the standard concentration result (e.g. [20, Theorem 1.2]) and an $\varepsilon$-net to cover $B$, we can see $\lim_{\mathbb{N}\to\infty}\delta_{N}=0$. Taking the expectation in the above display and choosing $\lambda=\delta_{N}^{\frac{1}{2}}$, we obtain (2.18).
@@
Let $f$ be the limit of $\mathcal{F}_{N}$ given by Lemma 2.5. The convexity of $\mathscr{P}^{h}$ proved in Lemma 2.2 gives the convexity of $f(0,\cdot)$. This allows us to represent $f$ in terms of the Hopf formula [2, 13]. We refer to [10, Theorem 3.13] for the version needed here. ∎

### v2

@@ Abstract.
Recently, [2] considered spin glass models with additional conventional order parameters characterizing single-replica properties. These parameters are distinct from the standard order parameter, the overlap, used to measure correlations between replicas. A “min-max” formula for the free energy was prescribed in [2]. We rigorously verify this prescription in the setting of vector spin glass models featuring additional deterministic spin interactions. Notably, our results can be viewed as a generalization of the Parisi formula for vector spin glass models in [44], where the order parameter for self-overlap is already present.
@@ 1. Introduction
It is well understood that the classical Sherrington–Kirkpatrick (SK) model has one order parameter that characterizes the correlations between replicas of the system, which is called the overlap [47]. In more general models, there can be other order parameters. For instance, in the vector spin glass, the self-overlap of a single replica comes into play. We call order parameters that characterize the properties of a single replica the conventional order parameter. In models where both types of order exist, [2] recently proposed a “min-max” prescription for the variational formula of free energy. We rigorously verify this in the setting of vector spin glass.
More precisely, we consider a Hamiltonian of the form $H_{N}(\sigma)+G_{N}(\sigma)$ at size $N$. Here, $H_{N}(\sigma)$ is the standard spin glass Hamiltonian with Gaussian disorder, and $G_{N}(\sigma)$ accounts for additional deterministic spin interactions. Also, $G_{N}$ can be very general so that $H_{N}(\sigma)+G_{N}(\sigma)$ cannot be rewritten as $H_{N}(\sigma)$ with shifted coupling coefficients. Let $\pi$ be the overlap parameter and $m$ be the conventional order parameter in this model. We show that the limit of free energy is given by
$\lim_{N\to\infty}\frac{1}{N}\mathbb{E}\log\int e^{H_{N}(\sigma)+G_{N}(\sigma)}\mathrm{d}\sigma={\color[rgb]{0,0,0}\sup_{m}\inf_{\pi}\mathcal{P}(\pi,m)}$
for some Parisi-type functional $\mathcal{P}$, verifying [2, Eq. (12)] (there, the free energy has an additional minus sign; so “min-max” there corresponds to “max-min” here). We also clarify that the quantity on the left-hand side is usually called the “pressure”, which differs from the “free energy” by a multiplicative factor of the temperature (absorbed into $H_{N}(\sigma)$ and kept implicit here). Here, for practical purposes, we call it “free energy” by slightly abusing the language.
@@
Our result allows for more general $m$. For instance, $m$ can be the parameter for the mean magnetization, moments of spins, the self-overlap, and combinations of these quantities. Regardless of the choice, $m$ has always to be optimized after the overlap $\pi$.
Our proof is based on interpolating between the free energy with Hamiltonian $H_{N}(\sigma)+G_{N}(\sigma)$ and the free energy with Hamiltonian $H_{N}(\sigma)-\frac{1}{2}\mathbb{E}H_{N}(\sigma)^{2}$ along a Hamilton–Jacobi equation. The simplest form of this technique exists for the Curie–Weiss (CW) model (see [26, Chapter 3]). Variants of this have been used to handle the self-overlap order parameter in [38, 18]. We will see that the free energy with the latter Hamiltonian is “pure” in the sense that the only order parameter is the overlap, the same as the SK model.
@@ 1.1. Setting
For $N\in\mathbb{N}$, we denote the spin configuration by $\sigma=(\sigma_{ki})_{1\leqslant k\leqslant D,\,1\leqslant i\leqslant N}$ which is a $D\times N$ matrix in $\mathbb{R}^{D\times N}$. We view each column vector $\sigma_{i}=(\sigma_{ki})_{1\leqslant k\leqslant D}$ in $\sigma$ as an $\mathbb{R}^{D}$-valued vector spin. We sample each vector spin independently from $P_{1}$ and denote the distribution of $\sigma$ by $P_{N}$. More precisely, we have $\mathrm{d}P_{N}(\sigma)=\otimes^{N}_{i=1}\mathrm{d}P_{1}(\sigma_{i})$.
@@
We call the term $\frac{N}{2}\xi\left(\frac{\sigma\sigma^{\intercal}}{N}\right)$ the self-overlap correction, which is is equal to $\frac{1}{2}\mathbb{E}H_{N}(\sigma)^{2}$ and resembles the drift term in an exponential martingale. It has been proved in [19, Corollary 8.3] that
@@
where the Parisi-type functional $\mathscr{P}$ is defined later in (2.3). Here, $\Pi$ is the collection of overlap parameters and there is no conventional order parameter.
@@
The additional term $NG(m_{N})$ is of the type in the generalized CW model. The natural order parameter arising from $NG(m_{N})$ characterizes the limit of $m_{N}$.
@@ 1.2. Main results
The free energy $F^{\mathsf{soc}}_{N}$ in (1.2) is “pure” since the overlap is the only order parameter in the formula. Further, we will show in Proposition A.1 of Appendix A that, with the self-overlap correction, all conventional orders are trivial in the sense that $m_{N}$ converges to a constant for any bounded $h$.
@@
In view of Corollary 1.2, without the self-overlap correction, it is necessary to include the self-overlap as part of the conventional order. From (1.12), it is also a minimal requirement. One can interpret this as that the self-overlap is the canonical conventional order parameter. This is not surprising because the self-overlap together with the overlap (which characterizes the correlation between different replicas) completely describes the entire overlap array.
Remark 1.6.
We comment on the convexity requirement (H4) on $\xi$. In [2], the authors stated below [2, Eq. (77)] that the convexity assumption is needed for technical reasons and that the min-max formula should hold without it. However, we disagree with this for the following reasons. It is argued in [35, Section 6] that the free energy in the bipartite model (where $\xi$ is not convex) does not admit a variational representation, or at least one that originates from rewriting the Parisi formula as a saddle point problem. Hence, it is not clear if, in general, there is a variational formula to begin with, when there is no additional conventional order.
@@ 1.3. Related works
As mentioned before, the source of motivation is from [2], which used the “min-max” prescription to clarify the relationship between quenched free energies, annealed free energies, and the overlap (which is called the “replica order parameter” therein). An earlier work that considered a more restrictive class of spin glass models with conventional order parameters is [32]. Since the self-overlap is one particular conventional order parameter, these connect with the author’s recent studies [18, 17] on the self-overlap in vector spin glasses. The idea of adding the self-overlap correction first appeared in the Hamilton–Jacobi approach to spin glass models [36, 38, 35, 37]. For more detail on this approach, we refer to [26].
Works on spin glasses with CW-type interactions include [25, 14]. The model in [25] has the mixed even-order SK Hamiltonian with additional term $\frac{\beta}{2N}\left(\sum_{i=1}^{N}\sigma_{i}\right)^{2}+\sum_{i=1}^{N}h_{i}\sigma_{i}$ with $\beta\geqslant 0$ and i.i.d. Gaussians $(h_{i})_{1\leqslant i\leqslant N}$. The variational formula for the free energy is obtained through a simple argument via usual approaches from the CW model. In [14], the authors considered the SK Hamiltonian plus $\frac{\nu}{2N}\left(\sum_{i=1}^{N}\xi_{i}\sigma_{i}\right)^{2}+\lambda\sum_{i=1}^{N}\xi_{i}\sigma_{i}$ with $\nu\geqslant 0$, $\lambda\in\mathbb{R}$, and i.i.d. random variables $(\xi_{i})_{1\leqslant i\leqslant N}$ with finite fourth moment. The proof relies on the adaptive interpolation method introduced in [3, 4]. The model in [25] can be covered here by absorbing the centered part of $\sum_{i=1}^{N}h_{i}\sigma_{i}$ into $H_{N}(\sigma)$ and we believe that the method demonstrated there does not work if the CW interaction is non-convex. While our results are not directly applicable to the model in [14], we believe that a modified argument similar to the one for statistical inference models in [15] could work.
Since the theme of this work is to find a formula for free energy, we review works along this direction. Parisi initially proposed the formula for the free energy in the SK model in [45, 46]. Guerra rigorously proved the upper bound in [29] and Talagrand proved the matching lower bound in [51]. Extensions were made to various settings: the SK model with soft spins [39], the scalar mixed $p$-spin model [41, 40], the multi-species model [7, 42], and the mixed $p$-spin model with vector spins [43, 44]. The Parisi formula for the balanced Potts spin glass was recently established in [13]. For spherical spins, Parisi-type variational formulas have been proved for the SK model [50], the mixed $p$-spin model [24], and the multi-species model [12].
As we are employing PDE tools in the study of spin glasses, we briefly describe the history of this approach. It was initiated by Guerra in [28] in the replica symmetric regime, which was later extended to the replica symmetry breaking scenario in [9] and explored further in various settings in [8, 1, 11, 10]. The application of the PDE techniques to the CW model starts in [6] and continues in [27]. Mathematically, the PDE approach was adopted by Mourrat in [36, 33]. Applications to statistical inference models also include [34, 16, 21, 15, 23]. The well-posedness of the PDEs arising from these contexts has been studied in [22, 20].
@@ 2. Proofs
The goal is to prove our main results. We start with some definitions and conditions on $\xi$. Let $\mathbf{S}^{D}_{+}$ denote the subset of $\mathbf{S}^{D}$ comprising of positive semi-definite matrices. A useful fact (see [30, Theorem 7.5.4]) is that if $a\in\mathbf{S}^{D}$, then
@@
The derivative $\nabla\xi:\mathbb{R}^{D\times D}\to\mathbb{R}^{D\times D}$ is defined with respect to the entry-wise inner product. This encompasses a broad class of models and we refer to [37, Section 6] for examples.
@@ 2.1. Parisi-type functional
$\Pi=\left\{\pi:[0,1]\to\mathbf{S}^{D}_{+}\ \big|\ \text{$\pi$ is left-continuous and increasing}\right\}$
@@
To define the Parisi functional, we recall the Ruelle probability cascade [49]. Let $\mathfrak{R}$ denote the Ruelle probability cascade with overlap uniformly distributed over $[0,1]$ (see [40, Theorem 2.17]). Precisely, $\mathfrak{R}$ is a random probability measure on the unit sphere in a separable Hilbert space, with the inner product denoted by $\alpha\wedge\alpha^{\prime}$. Independently sampling $\alpha$ and $\alpha^{\prime}$ from $\mathfrak{R}$, the law of $\alpha\wedge\alpha^{\prime}$ under $\mathbb{E}\mathfrak{R}^{\otimes 2}$ is the uniform distribution over $[0,1]$, where $\mathbb{E}$ integrates the randomness inherent in $\mathfrak{R}$. Almost surely, the support of $\mathfrak{R}$ is ultrametric in the induced topology. For rigorous definitions and comprehensive properties, we direct the reader to [40, Chapter2] (also see [26, Chapter5]).
@@
For the construction and properties of this process, we refer to [19, Section 4 and Remark 4.9].
@@
Here, $\mathbb{E}$ integrates the Gaussian randomness in $w^{\nabla\xi\circ\pi}(\alpha)$ and then the randomness of $\mathfrak{R}$. It is noteworthy that $\frac{1}{2}\nabla\xi\circ\pi(1)\cdot\tau\tau^{\intercal}=-\frac{1}{2}\mathbb{E}\left(w^{\nabla\xi\circ\pi}(\alpha)\cdot\tau\right)^{2}$, is the self-overlap correction in this functional.
@@ 2.2. Preliminaries
$\widetilde{\mathcal{F}}_{N}(t,x)=\frac{1}{N}\log\int\exp\left(H_{N}(\sigma)-\frac{N}{2}\xi\left(\frac{\sigma\sigma^{\intercal}}{N}\right)+tNG\left(m_{N}\right)+Nx\cdot m_{N}\right)\mathrm{d}P_{N}(\sigma).$
@@
The convergence at $x=0$ has been proved in [19, Corollary 8.3] (which is exactly (1.3)). For $x\neq 0$, we can substitute $\mathrm{d}P_{1}$ with $e^{x\cdot h(\tau)}\mathrm{d}P_{1}(\tau)$ and notice that the new measure still satisfies (H0). The convergence at $x$ follows from this substitution. ∎
@@
It is classical that (e.g. [48, Theorem 25.5]) the combination of convexity and differentiability implies continuous differentiability. Hence, it remains to show that $\mathscr{P}^{h}$ is differentiable everywhere. Fix any $x\in\mathbb{R}^{d}$. A vector $a\in\mathbb{R}^{d}$ is said to be a subdifferential of $\mathscr{P}^{h}$ at $x$ if $\mathscr{P}^{h}(y)-\mathscr{P}^{h}(x)\geqslant a\cdot(y-x)$ for every $y\in\mathbb{R}^{d}$. Since $\mathscr{P}^{h}$ is convex, it suffices to show that any subdifferential $a$ of $\mathscr{P}^{h}$ at $x$ is unique. For each $\varepsilon>0$, we choose $\pi_{\varepsilon}$ to satisfy
@@
  • for each $N$, $\mathcal{F}_{N}$ is Lipschitz and convex jointly over $\mathbb{R}_{+}\times\mathbb{R}^{d}$;
@@
For brevity, we write $m=m_{N}$ and let $\left\langle\cdot\right\rangle$ be the Gibbs measure naturally associated with $\mathcal{F}_{N}(t,x)$. The value of $(t,x)$, on which $\left\langle\cdot\right\rangle$ depends, will be clear from the context. At every $(t,x)\in(0,\infty)\times\mathbb{R}^{d}$, we can compute
@@ 2.3. A PDE approach
If $f$ is a limit of $\mathcal{F}_{N}$, then Lemma 2.1 implies that the relevant initial condition should be $f(0,\cdot)=\mathscr{P}^{h}$, which is continuously differentiable by Lemma 2.2. Also, Lemma 2.3 ensures that $f$ is Lipschitz and convex. These two properties make the following convex selection principle useful in this setting. We refer to [26, Theorem 3.21 and Corollary 3.24] for the proof. This result first appeared in [15].
@@
As $\to\infty, N\mathcal{F}_{N}$ converges in the local uniform topology to the unique viscosity solution $f$ of (2.13) with initial condition $f(0,\cdot)=\mathscr{P}^{h}$.
@@
Since $\mathcal{F}_{N}$ is Lipschitz uniformly in $N$ (by Lemma 2.3), the Arzelà–Ascoli theorem implies that any subsequence of $\mathcal{F}_{N}$ has a further subsequence that converges in the local uniform topology to some $f$. By the uniqueness of the solution (e.g. [26, Corollary 3.7]), it suffices to show that any such $f$ is the viscosity solution. For lighter notation, we assume that the entire sequence $\mathcal{F}_{N}$ converges to $f$.
@@
Using the convexity of $\widetilde{\mathcal{F}}_{N}$ due to (2.12), we have
$\widetilde{\mathcal{F}}_{N}(t_{N},x_{N}+y)\geqslant\widetilde{\mathcal{F}}_{N}(t_{N},x_{N})+y\cdot\nabla\widetilde{\mathcal{F}}_{N}(t_{N},x_{N}).$
@@
By the standard concentration result (e.g. [40, Theorem 1.2]) and an $\varepsilon$-net to cover $B$, we can see $\lim_{\mathbb{N}\to\infty}\delta_{N}=0$. Taking the expectation in the above display and choosing $\lambda=\delta_{N}^{\frac{1}{2}}$, we obtain (2.18).
@@
Let $f$ be the limit of $\mathcal{F}_{N}$ given by Lemma 2.5. The convexity of $\mathscr{P}^{h}$ proved in Lemma 2.2 gives the convexity of $f(0,\cdot)$. This allows us to represent $f$ in terms of the Hopf formula [5, 31]. We refer to [26, Theorem 3.13] for the version needed here. ∎